OpenStudy (anonymous):

Kim is showing her steps to solve the expression −18.9 ÷ 1 over 4. In which step did Kim make an error? Step 1: −18.9 ÷ 1 over 4 Step 2: −18.9 ⋅ 4 Step 3: 75.6

OpenStudy (anonymous):

step 1 step 2 step 3 or there is no error

OpenStudy (anonymous):

what do you think???

OpenStudy (igreen):

Is it \(\sf \dfrac{-18.9 \div 1}{4}\) or is it \(\sf -18.9 \div \dfrac{1}{4}\)

OpenStudy (igreen):

The 2nd one right? @alicelahey

OpenStudy (anonymous):

its the second step,

OpenStudy (igreen):

No..it's not

OpenStudy (igreen):

In the second step, she flipped 1/4 and multiplied..which gave her \(\sf -18.9 \cdot 4\)

OpenStudy (anonymous):

I meant the second one you put.

OpenStudy (igreen):

Oh, okay..

OpenStudy (igreen):

What's -18 * 4?

OpenStudy (igreen):

-18.9 * 4

OpenStudy (anonymous):

-75.6

OpenStudy (igreen):

Yes..so her problem was in Step 3, because she multiplied incorrectly.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.